not sure if in right spot but i need help bad
2. Clear fractions or decimals, solve, and check: 1/2+4m=3m-5/2
i dont understand what this means please help me.
does it mean this?
1/2 * 1/4 = 1/4
1/3 * -5/2 = 1/3 * -15/6 = -15/6
but i atill have fractions iam lost.

Answer by rfer(16322)

make all elements either fractions or decimals
I like decimals.
1/2+4m=3m-5/2
.5+4m=3m-2.5
subtract 3m
.5+m=-2.5
subtract .5
m=-2.5-.5
m=-3
------------------
or
1/2+8/2m=6/2m-5/2
2/2m=-6/2
m=-3